OpenStudy (anonymous):

given f(x)=1/x-4 abd g(x)=4-3x^2 find: g(-2) f o g (x) g(x^2+1) f^-1(x)

hero (hero):

g(x) = 4-3x^2 g(-2) = 4-3(-2)^2

hero (hero):

f o g(x) = f(g(x) = 1/((4-3x^2)-4)

hero (hero):

g(x^2-1) = 4-3(x^2-1)^2

hero (hero):

By the way, I haven't simplified any of those. You'll have to do that yourself. But at least I got you started

OpenStudy (anonymous):

yes I dont understand the second please explain

hero (hero):

lalaly, I see you

OpenStudy (lalaly):

lol everyone does ^_^

hero (hero):

Just press the I believe button: f o g(x) = f(g(x)) means that the entire equation that represents g(x) is a function of f(x). In other words, the whatever the equation for g(x) is, put that in the "x" portion of the equation for f(x), then simplify

hero (hero):

That's the best way I can describe it

OpenStudy (anonymous):

do I expand the brackects

hero (hero):

To simplify, yes

OpenStudy (anonymous):

complicated i can not

OpenStudy (anonymous):

can you help please